OpenStudy (anonymous):

Help.. on Differentials, see below problem. Use differentials to approximate the quantity. Sq root of 8.7

OpenStudy (anonymous):

\[\sqrt{x} + 1/2\sqrt{x}*Change\]

OpenStudy (anonymous):

So it is... \[\sqrt{9} + 1/(2\sqrt{9})*-.3\]

OpenStudy (amistre64):

y = sqrt(x); y' = 1/2sqrt(x) sqrt(8.7) is close to sqrt(9) y' = 1/2sqrt(9) = 1/6 linear approximation would be: a = 1/6 x -9/6 +3 a = (x+9)/6 ; when x = 8.7 we get a = (8.7+9)/6 = 17.7/6 = 2.95 2.95^2 = 8.7025

OpenStudy (anonymous):

hmmm, when I enter 8.7025 into my online hw it says its incorrect but doesn't tell me what is the correct answer. The answer must be 4 decimal places. This is what I did: f'(x) = (1/2) x^(-1/2) all mult by 0.3(which is the difference of 9 and 8.7) = 0.5

OpenStudy (anonymous):

It should be 2.95... Look above at my equation. It is f(x)+ Derivative of f(x) * the small amount

OpenStudy (amistre64):

 my idea, which prolly aint the right "method" is: find the equation of the tangent line near 8.7; say at 9 \[y=\frac{x+9}{6}\to\ \frac{17.7}{6}=2.95\] \[(2.95)^2=8.7025\]\[h=8.7025-8.7000=.0025\] \[\Delta y=hf'(9)=\frac{.0025}{6}\] drop the tangent equation by \(\Delta y\) \[y_n=\frac{x+9-.0025}{6}\] \[y_n=\frac{8.7+9-.0025}{6}\to\ \frac{17.6975}{6}=2.94958\bar 3\] \[(2.94958\bar 3)^2=8.7000418402\bar 7\]